OpenStudy (anonymous):

How do u solve by substitution 2x=3y x+2y=7

OpenStudy (anonymous):

2x=3y = x=3y/2, sub in for x in x+2y=7

OpenStudy (anonymous):

thanks

OpenStudy (gw2011):

2x=3y x=(3y)/2 Substitute into: x+2y=7 (3y)/2+2y=7 Multiply the above equation by 2 to get: 3y+4y=147y=14 y=2 Now substitute 2 for y in the equation: 2x=3y 2x=(3) (2)=6 x=3 Therefore, x=3 and y=2